OpenStudy (anonymous):

Can someone please help me with this question I'll give a medal if u explain and show me how

OpenStudy (anonymous):

The question is y > or greater than -x + 2 2 < or equal to x < or equal to 7 y < or equal to 1/2 + 5 f ( x, y) = 8x + 3y First it should be graphed then I should showmy shaded part. See if u could guide me through please

OpenStudy (anonymous):

Please anybody

OpenStudy (tkhunny):

\(y > -x + 2\) Graph this, please. It is already in slope-intercept form.

OpenStudy (anonymous):

Ok then wat?

OpenStudy (tkhunny):

Graph the next one. \(y \le (1/2)x + 5 \)

OpenStudy (anonymous):

What about 2 n 7?

OpenStudy (tkhunny):

Graph those, too. \( 2 \le x\) \( x \le 7 \) Get all four of them together and your feasible region should be well-defined.

OpenStudy (anonymous):

U mean one point is 2 by itself and then 7 by itself? !

OpenStudy (tkhunny):

It's not a point. It's a REGION. It extends vertically, both up and down.

OpenStudy (anonymous):

Umm.. can u explain more I dont kno what u mean

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

@Edge help me out in this question please

OpenStudy (anonymous):

ok im here

OpenStudy (anonymous):

I c that

OpenStudy (anonymous):

ok wait im doing it

OpenStudy (anonymous):

Done?